OpenStudy (anonymous):

Use the given function to compute the derivative. h(t) = 12√t + 6/√t

OpenStudy (anonymous):

h(t) = 12√t + 6/√t use the product rule on 12√t , and quotient rule on 6/√t

OpenStudy (anonymous):

d(h(t))/dt = 12d(√t)/dt + 6d(t)^-1/2

OpenStudy (anonymous):

h'(t)=12(1/2)t-1/2 +6(-1/2)t^-3/2 can you follow next?

OpenStudy (anonymous):

i just used the product rule

OpenStudy (anonymous):

Use the power rule, if you can, it's much simpler to work with.

OpenStudy (anonymous):

\[\Large h'(t) = {d\over dx}(12\sqrt{t}) \ + \ {d\over dx}(6\div\sqrt{t})\]

OpenStudy (anonymous):

\[\frac{ 12 }{ 2\sqrt{t} }+6 \times 2\sqrt{t}\]

OpenStudy (anonymous):

u Dont need Product Rule...To solve this...

OpenStudy (anonymous):

\[\sqrt{x} = x ^{\frac{ 1 }{ 2 }}\] \[\frac{ 6}{ \sqrt{t} } = 6*t ^{\frac{ -1 }{ 2 }}\]

OpenStudy (anonymous):

does my 12 dissappear?

OpenStudy (anonymous):

6+12t/t just wondering if that's right, and if its not just wondering what the right answer is so i can work thorough it.

OpenStudy (anonymous):

\[\frac{ 6 }{ \sqrt{t} } \frac{ -3 }{ \sqrt[\frac{ 3 }{ 2 }]{t} }\]

OpenStudy (anonymous):

is that the answer?